simplify
x+1
_____
x2-1x-2
: simplify
x+1
_____
x2-1x-2

Answer by jim_thompson5910(8196)


(x+1)/(x^2-x-2) Start with the given expression.


(x+1)/((x+1)(x-2)) Factor x^2-x-2 to get (x+1)(x-2).


(highlight(x+1))/(highlight(x+1)(x-2)) Highlight the common terms.


(cross(x+1))/(cross(x+1)(x-2)) Cancel out the common terms.


(1)/(x-2) Simplify.


So (x+1)/(x^2-x-2) simplifies to (1)/(x-2).


In other words, (x+1)/(x^2-x-2)=(1)/(x-2) where x<>-1 or x<>2
Question 150923: simplify
x+1
_____
x2-1x-2
: simplify
x+1
_____
x2-1x-2

Answer by jojo14344(326)

Start w/ the eqn:
=(x+1)/(x^2-x-2)
=cross(x+1)/(x-2)cross(x+1)
=1/(x-2)
=(x-2)^-1
Thank you,
Jojo